OpenStudy (ebonyserrano):

The sum of the four angle measures of any convex quadrilateral is 360 degrees. Suppose that a convex quadrilateral has angle measures of 60, 6y+10, and 8y-2 degrees. Write an expression for the degree measure of the fourth angle.

OpenStudy (ebonyserrano):

@Will.H

OpenStudy (will.h):

well so they want you to write an expressions for that fourth angle and the angle must measure 60 yet i think they gave us all the terms and they want us to find the fourth! hmm not so sure here but i will try 60 + 6y+10 + 8y - 2 = 360 14y +68 = 360 14y = 292 y = 20.857 bruh this looks confusing

OpenStudy (will.h):

wait wouldn't the fourth angle be just y! since it is unknown

OpenStudy (ebonyserrano):

no you did a great job, i understand the process of how you did

OpenStudy (ebonyserrano):

and yes

OpenStudy (will.h):

really so it is just y?

OpenStudy (will.h):

see if thats right

OpenStudy (ebonyserrano):

so what would the expression be?

OpenStudy (will.h):

we can't possibly figue out it is just (y)

OpenStudy (will.h):

 how can we determine that other angle expression!

OpenStudy (ebonyserrano):

it would be -14y+292

OpenStudy (will.h):

you can have it as (y) only that would make sense 60 + y + 8y -2 + 6y+10 = 360 15y +68 = 360 15y = 292 y = 19.4666666667

OpenStudy (will.h):

how did you know that